# LongForm: Effective Instruction Tuning with Reverse Instructions

### v2

@@ Abstract
Instruction tuning enables language models to more effectively generalize and better follow user intent. However, obtaining instruction data is costly and challenging. Prior work employs methods such as expensive human annotation, crowd-sourced datasets with alignment issues, and generating noisy examples via LLMs. We introduce the LongForm-C dataset, which is created by reverse instructions. We generate instructions via LLMs for human-written corpus examples using reverse instructions. First we select a diverse set of human-written documents from corpora such as C4 and Wikipedia; then we generate instructions for these documents via LLMs. This approach provides a cheaper and cleaner instruction-tuning dataset with natural output and one suitable for long text generation. Our models outperform 10x larger language models without instruction tuning on tasks such as story/recipe generation and long-form question answering. Moreover, LongForm models outperform prior instruction-tuned models such as FLAN-T5 and Alpaca by a large margin, and improve language understanding capabilities further. Finally, our models can effectively follow and answer multilingual instructions; we demonstrate this for news generation. We publicly release our data and models: https://github.com/akoksal/LongForm.
First published on ArXiv on 2023-04-17. This version extends the evaluation with new metrics and NLU tasks.
@@ 1 Introduction
Instruction tuning conditions language models (LMs) on user intents and improves cross-task generalization while ensuring better human alignment. Recent works on instruction tuning focus on data collection, showing that training with even a small amount of instruction-tuning data can outperform much larger LMs in following user intent and generalization Ouyang et al. (2022); Chung et al. (2022). However, these studies have some limitations, such as reliance on expensive human-annotated instruction data Ouyang et al. (2022) or focusing on instructions for academic NLP tasks that have limited coverage of real-world generation tasks Chung et al. (2022); Wang et al. (2022b); Sanh et al. (2022). A third approach, works that generate instructions and outputs from scratch using large language models (LLMs), often produces low-quality datasets Honovich et al. (2022); Wang et al. (2022b).
@@
Based on the LongForm-C dataset, we finetune instruction-following PLMs, which we call LongForm models, with different architectures and sizes: T5-XL Raffel et al. (2020), OPT-6.7B Zhang et al. (2022), and LLaMA-7B Touvron et al. (2023). We compare these models with the baselines FLAN-T5 Chung et al. (2022), T0pp Sanh et al. (2022), Tk-Instruct Wang et al. (2022b), and Alpaca on a diverse set of tasks: story, poem, email and recipe generation; grammar error correction; text summarization; table-to-text; and long form question answering. Our experimental results are as follows. (i) LongForm-OPT-2.7B outperforms OPT-30B on long text generation despite OPT-30B having 10x more parameters. (ii) LongForm models outperform prior instruction-following models, such as FLAN-T5 and Alpaca, with more than 63% (in-domain) and 35% (out-of-domain) relative METEOR improvement on text generation. (iii) The LongForm-C dataset enhances the language understanding capabilities of LLMs. It improves the relative performance of MMLU Hendrycks et al. (2021) by 10% when combined with FLAN. (iv) For multilingual news generation, LongForm models follow multilingual instructions and generate news articles in German, Spanish, French and Russian better than prior models.
@@
Figure 2: The reverse instructions method. After collecting diverse examples from corpora, we generate relevant instructions through zero-shot prompting via LLMs in various styles. (see §E.1)
@@ 2 Related Work
Instruction-following models: The instruction paradigm was introduced to better control language models using natural language commands Ouyang et al. (2022); Wang et al. (2022b); Chen et al. (2022); Wei et al. (2022). While some approaches Ouyang et al. (2022) focus on obtaining costly human-annotation data through their platform and improving human alignment with reinforcement learning, recent works Srivastava et al. (2022); Wang et al. (2022b) have extended this paradigm to various existing NLP tasks by reformulating them. This approach helps to overcome the instruction-data bottleneck, but is limited to academic tasks. Additionally, this approach is still restricted in terms of text generation: only 7 out of 1762 tasks in BigBench Srivastava et al. (2022) and 12 out of 1613 in Super-Natural Instructions (NIv2) Wang et al. (2022b) include English tasks with long outputs (i.e., average number of words $>$50). Self-Instruct Wang et al. (2022a) and Unnatural Instructions Honovich et al. (2022) challenge academic benchmarks by generating instruction data (input, instruction and output) via LLMs to provide a more diverse dataset. However, relying solely on generation via LLMs has drawbacks: the overall validity of Self-Instruct is only 54% Wang et al. (2022a).
@@
Data Generation with LMs: Recent neural network based approaches to NLP rely on large and diverse training datasets. Thus, many works focus on augmenting or generating new training examples. For example, existing datasets are expanded to enhance model quality for question answering (QA) Longpre et al. (2019), part-of-speech tagging Şahin and Steedman (2018), and textual similarity Schick and Schütze (2021). For a more general-purpose training, Self-Instruct Wang et al. (2022a) and Unnatural Instructions Honovich et al. (2022) propose the use of LLMs to generate tasks and examples including instructions and outputs. To the best of our knowledge, LongForm-C is the first work to combine corpora and LLMs to automatically generate a general-purpose text generation dataset. Corpus Mining: Most NLP tasks benefit from extracting useful examples from corpora such as machine translation (MT) with bitext mining Resnik (1999) and argument mining. Recent approaches use embeddings to find similar texts in different languages to mine human translation from unlabeled data Artetxe and Schwenk (2019). These methods have been extended to MT with up to 200 languages NLLB Team et al. (2022). Additionally, some tasks filter sentences from corpora to create diverse training sets for human annotation, such as argument mining Ein-Dor et al. (2020). Recent works explore methods to mine and restructure corpora (e.g., Wikipedia) to generate synthetic data for conversational QA Dai et al. (2022) and closed-book QA Lewis et al. (2021). Long Text Generation: Long text generation is challenging. It requires models to understand long dependencies and planning to generate long texts. Many works do task-specific generation (e.g., generation of stories Fan et al. (2018), recipes Bień et al. (2020), essays Feng et al. (2018) and long-form QA Fan et al. (2019)) by collecting useful examples from existing resources (e.g., ELI5 subreddit, WritingPrompts subreddit, recipe websites). Also, proper evaluation of long text generation remains a significant challenge for task-specific or general-purpose models Celikyilmaz et al. (2020).
@@
Reverse Instructions: The reverse instructions method offers a cost-effective and faster alternative for creating instruction tuning datasets compared to human annotation, while also yielding higher quality output compared to fully-synthetic data generation. Building upon this approach, Wang et al. (2023b) and Li et al. (2023) expand the concept of reverse instructions by integrating instruction data filtering mechanisms and open-source models. Furthermore, Chen et al. (2023) introduces a rewriting step of corpus samples in reverse instructions.
@@ 3 The LongForm-C Dataset
The LongForm-C dataset consists of 15,000 corpus examples generated via reverse instructions and additional 12,739 examples from structured corpora (SC) and NLP datasets with instructions and long text pairs; these two parts are referred to as SC and NLP below. The main part of the LongForm-C dataset is constructed by reverse instructions generation for a diverse set of corpus samples; this part is referred to as RI below. Additionally, we expand our dataset with SC examples through parsing and templates, as well as NLP tasks reformulated to increase diversity. Thus, our dataset features a diverse collection of instructions (including human-written, template-based and LLM-generated), each paired with a human-written output.
@@ 3.1 Reverse Instructions (RI)
For the English Wikipedia, we adopt a direct approach since it is already diverse. We randomly select 5,000 articles and extract the first paragraph for 75% and the first two paragraphs for 25% of the examples. The resulting dataset contains shorter texts: average length is 57$\pm$43 words, compared to 408$\pm$263 words for C4. The overall average in LongForm-C is 291$\pm$272 words. 2. Instruction Generation: We aim to generate relevant instructions for the 15K corpus examples using LLMs. We employ GPT3 (text-davinci-003 from OpenAI API) as our LLM. We design prompts for generating instructions for a given document and query the LLM in a zero-shot manner. We create three templates to diversify the styles of our instructions: formal instruction style (50%), informal chatbot style (30%), and search engine query style (20%). The formal instruction template is structured as follows: Instruction: X Output: “<corpus_example>” What kind of instruction could this be the answer to? X: We add “X:” at the end of the template as a prompt to generate plausible instructions. See §E.1 for details of LLM parameters and templates.
Finally, we incorporate length information into the generated instruction using a predefined set of templates to signal the desired length of the output and provide additional control to the model. As the length of our corpus examples varies (291$\pm$272 words), we provide templates such as “Respond in D sentences” or “Respond in D words”. We also include more ambiguous templates such as “Respond briefly” for outputs of less than 3 sentences, and “Respond in detail” for outputs of more than 10 sentences. We append or prepend these templates to the original instructions for a 30% subset of our dataset. This allows us to additionally control LongForm models in terms of length.
@@ 3.2 Structured Corpora (SC)
We have focused on reverse instructions generation so far. But there are also structured corpora that contain instruction-output pairs. We collect data from Stack Exchange and WikiHow without using LLMs. Stack Exchange (SE): We use the Stack Exchange subcorpus of Pile Gao et al. (2020). We select 50 examples from each of 88 subdomains, covering a wide range of topics, including Buddhism, chemistry, and webmasters. For each example, we select the question and its corresponding details as an instruction and the answer as the output. This subset adds more complicated human instructions to LongForm-C. WikiHow: WikiHow includes how-to tutorials, each comprising a question and an answer that consists of two parts: an introduction and several steps with a brief summary and a long description of each step.
To generate an instruction, we create 18 templates – e.g., “What are some steps to <question>?” – and use “how-to” questions to fill in the placeholders. We also include the number of steps in 14 out of 18 templates to control the model in terms of number of steps, similar to the length information for corpus-extracted examples. We provide all templates in §E.2.
We generate target texts by combining various elements, including introductory paragraphs (i.e., part 1), summaries and long descriptions of each step (i.e., part 2). We only include the first part in half of the examples. The second part is always included, but in different formats: either as a summary, a long description, or both. The final size of our WikiHow subset is 2,500 examples.
@@ 3.3 NLP Tasks (NLP)
Most of instruction tuning benchmarks of NLP tasks involve classification or short text generation tasks, accounting for over 99% of BigBench and NIv2. However, we opt for NLP tasks with long outputs to enrich LongForm-C. From BigBench, we select Helpful Answer Generation (hhh) Askell et al. (2021) and Minute Mysteries QA, a question answering dataset for short stories. From NIv2, we include a selection of ten tasks including poem generation Hipson and Mohammad (2020), table to natural text generation Iyyer et al. (2017), story generation Sap et al. (2020); Orbach and Goldberg (2020); Lin et al. (2019), text summarization Fabbri et al. (2019); Kornilova and Eidelman (2019); Bražinskas et al. (2020), and fact generation Kotonya and Toni (2020). For a more uniform distribution across datasets, we sample a similar number from each dataset, resulting in 600 samples from BigBench and 3,684 from NIv2. Finally, we include the Enron email dataset to create an email writing task from the subject using templates similar to WikiHow. We also employ the BEA-2019 shared task on grammatical error correction Bryant et al. (2019) by using incorrect inputs and template instruction as the input and the correct output as the target. We provide templates for Enron in §E.3 and for BEA-2019 in §E.4.
3.4 Dataset Analysis
Diversity: To analyze the diversity of the generated instructions, we adopt the approach proposed by Wang et al. (2022a). We parse each instruction in the corpus subset with Berkeley Neural Parser Kitaev and Klein (2018). We extract either the verb and its direct object (noun+verb: “describe history”) or the auxiliary with its dependent verb (auxiliary+verb: “do know”). 8,872 of 15,000 instructions follow this format. Figure 3 visualizes the 3,167 most common noun+verb/auxiliary+verb pairs. They cover a diverse range of topics, including events, careers, biographies, and writing tasks such as reviews, summaries and essays – and this is just looking at the 11% of our LongForm-C dataset with reverse instructions (RI).
@@ 3.4 Dataset Analysis
Table 1: Origin statistics for LongForm-C. Examples are generated with reverse instructions, from structured corpora (SC) and NLP tasks (NLP) such as email writing, grammar error correction, story/poem generation and text summarization.

### v3

# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt

@@ Abstract
Instruction tuning enables language models to more effectively generalize and better follow user intent. However, obtaining instruction data is costly and challenging. Prior work employs methods such as expensive human annotation, crowd-sourced datasets with alignment issues, and generating noisy examples via LLMs. We introduce the LongForm-C dataset, which is created by reverse instructions. We generate instructions via LLMs for human-written corpus examples using reverse instructions. First we select a diverse set of human-written documents from corpora such as C4 and Wikipedia; then we generate instructions for these documents via LLMs. This approach provides a cheaper and cleaner instruction-tuning dataset with natural output and one suitable for long text generation. Our models outperform 10x larger language models without instruction tuning on tasks such as story/recipe generation and long-form question answering. Moreover, LongForm models outperform prior instruction-tuned models such as FLAN-T5 and Alpaca by a large margin, and improve language understanding capabilities further. We publicly release our data and models: https://github.com/akoksal/LongForm.
First published on ArXiv on 2023-04-17. This version extends the training with recent LLMs, evaluation with new metrics, and NLU tasks.
@@ 1 Introduction
Instruction tuning conditions language models (LMs) on user intents and improves cross-task generalization while ensuring better human alignment. Recent works on instruction tuning focus on data collection, showing that training with even a small amount of instruction-tuning data can outperform much larger LMs in following user intent and generalization Ouyang et al. (2022); Chung et al. (2024). However, these studies have some limitations, such as reliance on expensive human-annotated instruction data Ouyang et al. (2022) or focusing on instructions for academic NLP tasks that have limited coverage of real-world generation tasks Chung et al. (2024); Wang et al. (2022); Sanh et al. (2022). A third approach, works that generate instructions and outputs from scratch using large language models (LLMs), often produces low-quality datasets Honovich et al. (2023); Wang et al. (2022).
@@
Based on the LongForm-C dataset, we finetune instruction-following PLMs, which we call LongForm models, with different architectures and sizes: T5-XL Raffel et al. (2020), OPT-6.7B Zhang et al. (2022), and LLaMA-7B Touvron et al. (2023a). We compare these models with the baselines FLAN-T5 Chung et al. (2024), T0pp Sanh et al. (2022), Tk-Instruct Wang et al. (2022), and Alpaca on a diverse set of tasks: story, poem, email and recipe generation; grammar error correction; text summarization; table-to-text; and long form question answering. Our experimental results are as follows. (i) LongForm-OPT-2.7B outperforms OPT-30B on long text generation despite OPT-30B having 10x more parameters. (ii) LongForm models outperform prior instruction-following models, such as FLAN-T5 and Alpaca, with more than 63% (in-domain) and 35% (out-of-domain) relative METEOR improvement on text generation. (iii) In addition to NLG, the LongForm-C dataset also enhances the language understanding (NLU) capabilities of several LLMs. It improves performance on MMLU Hendrycks et al. (2021) compared to prior datasets like FLAN.
(iv) For multilingual news generation, LongForm models follow multilingual instructions and generate news articles in German, Spanish, French and Russian better than prior models.
@@ 2 Related Work
Instruction-following models: The instruction paradigm was introduced to better control language models using natural language commands Ouyang et al. (2022); Wang et al. (2022); Chen et al. (2022); Wei et al. (2022). While some approaches Ouyang et al. (2022) focus on obtaining costly human-annotation data through their platform and improving human alignment with reinforcement learning, recent works Srivastava et al. (2023); Wang et al. (2022) have extended this paradigm to various existing NLP tasks by reformulating them. This approach helps to overcome the instruction-data bottleneck, but is limited to academic tasks. Additionally, this approach is still restricted in terms of text generation: only 7 out of 1762 tasks in BigBench Srivastava et al. (2023) and 12 out of 1613 in Super-Natural Instructions (NIv2) Wang et al. (2022) include English tasks with long outputs (i.e., average number of words $>$50). Self-Instruct Wang et al. (2023b) and Unnatural Instructions Honovich et al. (2023) challenge academic benchmarks by generating instruction data (input, instruction and output) via LLMs to provide a more diverse dataset. However, relying solely on generation via LLMs has drawbacks: the overall validity of Self-Instruct is only 54% Wang et al. (2023b). More recent works focus on curating costly human-annotated datasets such as Dolly Conover et al. (2023).
@@
Data Generation with LMs: Recent neural network based approaches to NLP rely on large and diverse training datasets. Thus, many works focus on augmenting or generating new training examples. For example, existing datasets are expanded to enhance model quality for question answering (QA) Longpre et al. (2019), part-of-speech tagging Şahin and Steedman (2018), and textual similarity Schick and Schütze (2021). For a more general-purpose training, Self-Instruct Wang et al. (2023b) and Unnatural Instructions Honovich et al. (2023) propose the use of LLMs to generate tasks and examples including instructions and outputs. WizardLM Xu et al. (2024) focuses on evolving human-written instructions by prompting LLMs and generating more complicated instructions that require multiple-step reasoning. To the best of our knowledge, LongForm-C is the first work to combine corpora and LLMs to automatically generate a general-purpose text generation dataset.
Corpus Mining: Most NLP tasks benefit from extracting useful examples from corpora such as machine translation (MT) with bitext mining Resnik (1999) and argument mining. Recent approaches use embeddings to find similar texts in different languages to mine human translation from unlabeled data Artetxe and Schwenk (2019). These methods have been extended to MT with up to 200 languages NLLB Team et al. (2022). Additionally, some tasks filter sentences from corpora to create diverse training sets for human annotation, such as argument mining Ein-Dor et al. (2020). Recent works explore methods to mine and restructure corpora (e.g., Wikipedia) to generate synthetic data for conversational QA Dai et al. (2022) and closed-book QA Lewis et al. (2021).
Long Text Generation: Long text generation is challenging. It requires models to understand long dependencies and planning to generate long texts. Many works do task-specific generation (e.g., generation of stories Fan et al. (2018), recipes Bień et al. (2020), essays Feng et al. (2018) and long-form QA Fan et al. (2019)) by collecting useful examples from existing resources (e.g., ELI5 subreddit, WritingPrompts subreddit, recipe websites). Also, proper evaluation of long text generation remains a significant challenge for task-specific or general-purpose models Celikyilmaz et al. (2020).
@@
Reverse Instructions: The reverse instructions method offers a cost-effective and faster alternative for creating instruction tuning datasets compared to human annotation, while also yielding higher quality output compared to fully-synthetic data generation. Building upon this approach, Wang et al. (2023c) and Li et al. (2023) expand the concept of reverse instructions by integrating instruction data filtering mechanisms and open-source models. Furthermore, Chen et al. (2024) introduces a rewriting step of corpus samples in reverse instructions. In a significant extension to multilingual settings, Köksal et al. (2024) proposed MURI (Multilingual Reverse Instructions), which leverages multilingual corpora and translation models to generate instruction-following datasets for 200 languages. Moreover, Ziegler et al. (2024) introduced CRAFT, which extends the reverse instructions methodology to task-specific contexts through the generation of structured datasets via corpus example retrieval and LLM-based reformulation.
The reverse instructions methodology constitutes an efficacious and cost-efficient alternative to conventional human annotation for instruction tuning dataset creation, while demonstrating superior quality compared to fully-synthetic data generation approaches. This framework has been subsequently extended through several significant contributions in the literature. Specifically, Wang et al. (2023c) and Li et al. (2023) augmented the fundamental reverse instructions paradigm through the integration of instruction data filtering mechanisms and the utilization of open-source models. Chen et al. (2024) further advanced this methodology by introducing a corpus sample rewriting procedure within the reverse instructions framework. Köksal et al. (2024) expanded it to multilingual settings via MURI (Multilingual Reverse Instructions), which leverages multilingual corpora with translation models and LLMs to generate instruction-following datasets for 200 languages. Furthermore, Ziegler et al. (2024) developed CRAFT, adapting the reverse instructions approach to generate task-specific structured datasets through corpus example retrieval and reformulation with LLMs.
Figure 2: The reverse instructions method. After collecting diverse examples from corpora, we generate relevant instructions through zero-shot prompting via LLMs in various styles. (see §F.1)
@@ 3 The LongForm-C Dataset
The LongForm-C dataset consists of 15,000 corpus examples generated via reverse instructions and additional 12,739 examples from structured corpora (SC) and NLP datasets with instructions and long text pairs. The main part of the LongForm-C dataset is constructed by reverse instructions generation for a diverse set of corpus samples; this part is referred to as RI below. Additionally, we expand our dataset with structured corpora examples through parsing and templates (referred to as SC), as well as NLP tasks reformulated to increase diversity (referred to as NLP). Thus, our dataset features a diverse collection of instructions (including human-written, template-based and LLM-generated), each paired with a human-written output.
@@ 3.1 Reverse Instructions (RI)
For the English Wikipedia, we adopt a direct approach since it is already diverse. We randomly select 5,000 articles and extract the first paragraph for 75% and the first two paragraphs for 25% of the examples. The resulting dataset contains shorter texts: average length is 57$\pm$43 words (compared to 408$\pm$263 for C4). The overall average in LongForm-C is 291$\pm$272.
2. Instruction Generation: We aim to generate relevant instructions for the 15K corpus examples using LLMs. We employ GPT3 (text-davinci-003 from OpenAI API) as our LLM. We design prompts for generating instructions for a given document and query the LLM in a zero-shot manner. We create three templates to diversify the styles of our instructions: formal instruction style (50%), informal chatbot style (30%), and search engine query style (20%). The formal instruction template is structured as follows: {mdframed}[style=MyFrame,hidealllines=true,backgroundcolor=gray!5] Instruction: X
Output: “<corpus_example>”
What kind of instruction could this be the answer to?
X: We add “X:” at the end of the template as a prompt to generate plausible instructions. See §F.1 for details of LLM parameters and templates.
Finally, we incorporate length information into the generated instruction using a predefined set of templates to signal the desired length of the output and provide additional control to the model. As the length of our corpus examples varies (291$\pm$272 words), we provide templates such as “Respond in D sentences” or “Respond in D words”. We also include less precise templates such as “Respond briefly” (resp. “Respond in detail”) for outputs of less than 3 (resp. more than 10) sentences. We append or prepend these templates to the original instructions for a 30% subset of our dataset. This lets us control LongForm models in terms of length.
@@ 3.2 Structured Corpora (SC)
We also mine structured corpora that contain (instruction,(long-)output) pairs: Stack Exchange and WikiHow. We do not use LLMs here.
From the Stack Exchange (SE) subcorpus of Pile Gao et al. (2020), we select 50 examples from each of 88 subdomains, covering a wide range of topics, including Buddhism, chemistry, and webmasters. For each example, we select the question and its corresponding details as an instruction and the answer as the output. This subset adds more complicated human instructions to LongForm-C.
WikiHow includes how-to tutorials, each comprising a question and an answer that consists of two parts: an introduction and several steps with a brief summary and a long description of each step.
To generate an instruction, we create 18 templates – e.g., “What are some steps to <question>?” – and use “how-to” questions to fill in the placeholders. We also include the number of steps in 14 out of 18 templates to control the model in terms of number of steps, similar to the length information for corpus-extracted examples. We provide all templates in §F.2.
We generate target texts by combining various elements, including introductory paragraphs (i.e., part 1), summaries and long descriptions of each step (i.e., part 2). We only include the first part in half of the examples. The second part is always included, but in different formats: either as a summary, a long description, or both.
@@ 3.3 NLP Tasks (NLP)
Most of instruction tuning benchmarks of NLP tasks involve classification or short text generation tasks, accounting for over 99% of BigBench and NIv2. However, we opt for NLP tasks with long outputs to enrich LongForm-C. From BigBench, we select Helpful Answer Generation (hhh) Askell et al. (2021) and Minute Mysteries QA, a question answering dataset for short stories. From NIv2, we include a selection of ten tasks including poem generation Hipson and Mohammad (2020), table to text generation Iyyer et al. (2017), story generation Sap et al. (2020); Orbach and Goldberg (2020); Lin et al. (2019), summarization Fabbri et al. (2019); Kornilova and Eidelman (2019); Bražinskas et al. (2020), and fact generation Kotonya and Toni (2020). For a more uniform distribution across datasets, we sample a similar number from each dataset, resulting in 600 samples from BigBench and 3,684 from NIv2.
Finally, we include the Enron dataset to create an email writing task from the subject using templates similar to WikiHow. We also employ the BEA-2019 for grammatical error correction Bryant et al. (2019); incorrect inputs and template instruction are the input, the correct output the target. See §F.3 for templates for Enron and §F.4 for BEA-2019.
@@
Table 1: Origin statistics for LongForm-C. Examples are generated with reverse instructions, from structured corpora (SC) and NLP tasks (NLP).
3.4 Dataset Analysis
Diversity: Following Wang et al. (2023b), we analyze the diversity of the generated instructions by parsing each instruction in the corpus subset with Berkeley Neural Parser Kitaev and Klein (2018). We extract either the verb and its direct object (noun+verb: “describe history”) or the auxiliary with its dependent verb (auxiliary+verb: “do know”). 8,872 of 15,000 instructions follow this format. Figure 3 visualizes the 3,167 most common noun+verb/auxiliary+verb pairs. They cover a diverse range of topics, including events, careers, biographies, and writing tasks such as reviews, summaries and essays – and this is just looking at the 11% of our LongForm-C dataset with reverse instructions (RI). For a detailed comparison of LongForm-C’s instruction diversity with other datasets such as Dolly (Conover et al., 2023) and FLAN (Chung et al., 2024), see Appendix (§G).